### v2

# L’approximation arithmétique de Fujita sur courbes adéliques

Abstract
In this paper, we will prove an analogue of Fujita’s approximation under the framework of Arakelov theory over adelic curves, which gives a positive response to a conjecture of Huayi Chen and Atsushi Moriwaki.
{altabstract}
Dans cet article, on démontrera une analogie de l’approximation de Fujita sous le cadre de la théorie d’Arakelov sur courbes adéliques, qui donne une réponse positive d’une conjecture de Huayi Chen et Atsushi Moriwaki.
1 Introduction
Fujita approximation is an approximative version of Zariski decomposition of pseudo-effective divisors, and see [28] for more details. More precisely, let $X$ be a projective variety over a field $k$, and $L$ be a line bundle on $X$. The volume of $L$ is defined as
$\operatorname{vol}(L)=\limsup_{D\rightarrow\infty}\frac{\dim_{k}H^{0}(X,L^{\otimes D})}{D^{\dim(X)}/\dim(X)!}.$
We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$. Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed as the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
$p^{-\dim(X)}\operatorname{vol}(A)\geqslant\operatorname{vol}(L)-\epsilon.$
Intuitively, this result asserts that a big line bundle can be arbitrarily closely approximated by ample line bundles along an effective direction.
This approximated decomposition theorem is proved by T. Fujita [12] for the case of characteristic 00, and by S. Takagi [24] for an arbitrary characteristic. We refer the readers to [17, §11.4] for a survey on this topic.
1.1 Arithmetic analogue over number fields
It is quite nature to ask whether there is an analogue of Fujita’s approximation theorem under the framework of Arakelov geometry, since we can define the arithmetic positivity and arithmetic volume of an adelic line bundle over arithmetic varieties.
1.1.1
The arithmetic analogue of volume function and arithmetic bigness under the framework of Arakelov geometry are introduced in [19, 20] by A. Moriwaki. Soon after these, voluminous works appeared on this topic, such as [21, 6, 27, 14, 22, 15].
Let $K$ be a number field, $\mathscr{X}$ be an arithmetic variety over $\operatorname{Spec}\mathcal{O}_{K}$ of dimension $d$, and $\overline{\mathscr{L}}$ be a Hermitian line bundle on $\mathscr{X}$. We define
$\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})=\log\#\left\{s\in H^{0}(\mathscr{X},\mathscr{L}^{\otimes D})\mid\|s\|_{\sup,v}\leqslant 1,\;\forall\;v\in M_{K,\infty}\right\},$
and
$\widehat{\operatorname{vol}}(\overline{\mathscr{L}})=\limsup_{n\to\infty}\frac{\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes n})}{n^{\dim(\mathscr{X})}/\dim(\mathscr{X})!}.$
By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar properties to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$, so it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which means $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the property of Fujita approximation. In fact, this used to be a conjecture of A. Moriwaki.
1.1.2
In [4, 26], H. Chen and X. Yuan proved the arithmetic Fujita approximation over number field at almost the same time. Although they both applied [18, §3], but their methods are essential different. X. Yuan [26, Theorem C] takes pluri-subharmonic envelope to deal with the infinite part, which is well-known in complex analysis.
In [4], H. Chen applied his $\mathbb{R}$-filtration method developed in [5]. More precisely, by considering the successive minima on $H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$, we can make $\bigoplus\limits_{n\geqslant 0}H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$ to be a filtrated graded linear series. Then by combining some results of approximable graded linear series in [18], we are able to apply some convergent properties of $\mathbb{R}$-filtrations to construct an ample line bundle which approximates $\overline{\mathscr{L}}$ well. This is the main idea of the proof in [4].
1.2 Arakelov geometry over adelic curves
In [9], H. Chen and A. Moriwaki developed a kind of Arakelov theory over the so-called adelic curve. Based on this framework, they built an arithmetic intersection theory in [10], and studied the arithmetic positivities in [11]. Under this setting, we can study the arithmetic geometry over number fields, function fields and other bases with adelic structures under a same framework.
1.2.1
Briefly speaking, an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ consists of an arbitrary field $K$ which is equipped with a parameter set $\Omega$ of all its absolute values. We require the set $\Omega$ parameterized by $\phi$ has a $\sigma$-algebra structure $\mathcal{A}$ and a measure $\nu$ compatible with $\mathcal{A}$.
Under this setting, we can generalize many classic results and subjects of Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula. Since the adelic structure becomes more complicated than those of classic cases, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique. In fact, some classic properties are not valid any longer in this framework, though we have developed many in [9, 10, 11], and see also [23].
1.2.2
Inspired by [4, 26], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves. Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina directly in [4] At the same time, some constructions do not work any longer without the help of integral models.
In [11, §8.5], H. Chen and A. Moriwaki proved a similar result to the arithmetic Fujita approximation over adelic curves, which is called the relative Fujita approximation. More precisely, let $\overline{L}$ be an adelic line bundle, where $L$ is big on $X$. Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a sufficiently large $p\in\mathbb{N}$ and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}\overline{L})^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$. Here $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ denote the asymptotic maximal and minimal slopes respectively.
Compared with this relative version above, the classic arithmetic Fujita approximation seems to take an "average" at every minima of minimal slope, and approximates all of them simultaneously, not only the maximal one. In fact, this is conjectured by Chen-Moriwaki in [11, Remark 9.2.7].
1.3 Main result
In this article, we prove the following result on the arithmetic Fujita approximation over adelic curves under the framework [9, 11], which gives a positive response to the conjecture [11, Remark 9.2.7]. More generally, we prove a version of adelic $\mathbb{R}$-line bundle below at the same time, which is also a generalization of [4, Theorem 4.3] and [26, Theorem C].
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, which satisfies either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$ over $S$. Then for any $\epsilon>0$, there exist a sufficiently large $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle. Moreover, we have
$p^{-(\dim(X)+1)}\widehat{\operatorname{vol}}(\overline{A})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon.$
1.3.1
In [11, §3], they develop a method which constructs a graded linear series over the adelic curve of trivially valued field from general one by its Harder-Narasimhan filtration. Next, we can modify the original norms by its spectral norm, which makes the graded linear series to be filtrated. Under this construction, the new graded linear series has some asymptotic invariants same as the original one. Then some ideas in [4, 18] can be absorbed in this work.
Once we prove the case of $\mathbb{Q}$-line bundle (in Corollary 5.2.1) from the case of usual adelic line bundle (in Corollary 5.1), we can approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle in an effective direction, and then we can prove the case of $\mathbb{R}$-line bundle (in Theorem 5.2.2).
1.3.2
In the application of $\mathbb{R}$-filtration method, we need to take some limits of some graded linear series and measures, and most operations only work when the objects are measurable and integrable. If either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, there will be no obstructions on the measurability or integrability, and see [9, §4] for more details of the necessity.
In order to study some arithmetic positivities by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$, we need to suppose the base field $K$ is perfect. For example, for an adelic line bundle $\overline{L}$ where $L$ is nef, we need to apply $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})\geqslant 0$ to deduce $\overline{L}$ is arithmetically nef by [11, Proposition 9.1.7], where we need the perfect condition. See [11, §9] for more details of the necessity.
Of course, in [4, 26], there is no such problem since for the case of number field, because $\mathcal{A}$ is discrete in the language of adelic curve and the base field is perfect by its characteristic 00.
1.4 Organization of article
In §2, we introduce the required notions of Arakelov theory over adelic curves and $\mathbb{R}$-filtration. In §3, we consider graded linear series over a general adelic curve and construct those over the adelic curve of trivially valued field filtrated by the Harder-Narasimhan filtration. In §4, we prove the property of Fujita approximation of certain graded linear series. In §5, we prove the arithmetic Fujita approximation of adelic line bundles in the usual case, the case of $\mathbb{Q}$-line bundle and the case of $\mathbb{R}$-line bundle sequentially.
Acknowledgements
I would like to thank Prof. Atsushi Moriwaki for introducing this subject to me. In addition, I would like to thank Prof. Huayi Chen for many useful discussions and suggestions, and Dr. Wenbin Luo for the advices on some key techniques. Part of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous invitation and reception. Chunhui Liu was supported by National Science Foundation of China No. 12201153.
2 Arakelov theory over adelic curves
In this section, we will introduce the foundations of Arakelov theory over adelic curves and of $\mathbb{R}$-filtration. The main references are [9, 11].
2.1 Foundations of Arakelov geometry over adelic curves
First, we recall some useful knowledge on the Arakelov geometry over adelic curves, and see [9] for a systematical introduction to this subject.
2.1.1 Adelic curve
Let $K$ be a field. We denote by $M_{K}$ the set of all absolute values on $K$.
[…]
We say that $(\Omega,\mathcal{A},\nu)$ is an adelic structure on $K$ equipped with a mapping
$\begin{array}[]{rrcl}\phi:&\Omega&\longrightarrow&M_{K}\\ &\omega&\mapsto&|\raisebox{1.72218pt}{.}|_{\omega}\end{array}$
which satisfies the properties below:
  1. (1)
    $\mathcal{A}$ is a $\sigma$-algebra on $\Omega$, and $\nu$ is a measure on $(\Omega,\mathcal{A})$;
  2. (2)
    for all $a\in K^{\times}$, the function
    $\begin{array}[]{rcl}\Omega&\longrightarrow&\mathbb{R}\\ w&\mapsto&\log|a|_{w}\end{array}$
    is $\mathcal{A}$-measurable and integrable with respect to $\nu$.
In addition, the data $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is called an adelic curve.
We denote by $\Omega_{f}$ the subset of $\Omega$ whose images in $M_{K}$ with respect to $\phi$ is non-archimedean, and by $\Omega_{\infty}$ the subset of $\Omega$ whose image in $M_{K}$ is archimedean. We have $\Omega=\Omega_{f}\sqcup\Omega_{\infty}$ by definition directly.
[…]
We say that an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is proper if for all $a\in K^{\times}$, we always have
$\int_{\Omega}\log|a|_{w}\nu(\mathrm{d}w)=0,$
and the above equality is called the product formula of $S$.
We refer the readers to [9, §3] for more details on these notions.
[…]
Let $K$ be a number field or a function field. Then the usual adèle ring of $K$ gives an adelic structure, and constructs examples of proper adelic curves.
Let $K$ be a finitely transcendental extension over $\mathbb{Q}$, which is a sub-field of $\mathbb{C}$. In [19], A. Moriwaki constructed a special adelic structure over $K$, and defined a height function with respect to it. This height function has been applied to study many arithmetic geometry subjects.
[…]
Let $K$ be an arbitrary field, and |.$|\raisebox{1.72218pt}{.}|_{0}$ be the trivial absolute value over $K$, which means
$|a|_{0}=\begin{cases}1,\text{ if }a\neq 0,\\ 0,\text{ if }a=0.\end{cases}$
We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ once. By definition, $S_{0}$ is a proper adelic curve.
Example 2.1.1 will be used in the article later. We refer the readers to [9, §3.2] for more examples of adelic curve.
2.1.2 Adelic vector bundle
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $E$ be a vector space over $K$. For a family of norms (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E_{w}=E\otimes_{K}K_{w}$ over $K_{w}$, if the norm family (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E$ and its dual norm family (.$(\|\raisebox{1.72218pt}{.}\|^{\vee}_{w})_{w\in\Omega}$ on $E^{\vee}$ are both $\mathcal{A}$-measurable on $\Omega$, then we say $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ is an adelic vector bundle over $S$. If $\dim(E)=1$, we say that $\overline{E}$ is an adelic line bundle over $S$.
We say that an adelic vector bundle $\overline{F}=(F,(\|\raisebox{1.72218pt}{.}\|^{\prime}_{w})_{w\in\Omega})$ is an adelic sub-bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $F$ is a $K$-vector subspace of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime}_{w}$ is the restriction of .$\|\raisebox{1.72218pt}{.}\|_{w}$ on $F$ for each $w\in\Omega$.
An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical quotient homomorphism $E\rightarrow G$ for each $w\in\Omega$. If $G$ is the quotient space of $E$ modulo $F$, and is equipped with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the $K$-vector space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E)_{w}$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$. We refer the readers to [9, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
[…]
In general, for an adelic vector bundle $\overline{E}$ over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, we cannot make sure that all its subspace equipped with the subspace norms is still an adelic vector bundle, and its quotient spaces admit the same situation. However, if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, we will get adelic vector bundles from all the above constructions, and see [9, Proposition 4.1.24] for more details. In the whole article, we will always assume these conditions.
2.1.3 Some Arakelov invariants
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, and $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ be an adelic vector bundle over $S$. For the determinant bundle $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ of $\overline{E}$, we select an $s\in\det(E)\smallsetminus\{0\}$, and define
$\widehat{\deg}(\overline{E})=-\int_{\Omega}\log\|s\|_{\det,w}\nu(\mathrm{d}w),$
which is well-defined due to [9, §4.2]. This is independent of the choice of $s$ by the product formula of $S$.
We define the positive Arakelov degree of $\overline{E}$ as
$\widehat{\deg}_{+}(\overline{E})=\sup_{F\subseteq E}\widehat{\deg}(\overline{F}),$
where $\overline{F}$ runs over all the adelic sub-bundles of $\overline{E}$.
For a non-zero adelic bundle $\overline{E}$, we define the slope of $\overline{E}$ as
$\widehat{\mu}(\overline{E})=\frac{\widehat{\deg}(\overline{E})}{\dim(E)}.$
In addition, we define the maximal slope of $\overline{E}$ as
$\widehat{\mu}_{\max}(\overline{E})=\sup_{F\subseteq E}\widehat{\mu}(\overline{F}),$
where $\overline{F}$ runs over all adelic sub-bundles of $\overline{E}$. Similarly, we define the minimal slope of $\overline{E}$ as
$\widehat{\mu}_{\min}(\overline{E})=\inf_{E\twoheadrightarrow G}\widehat{\mu}(\overline{G}),$
where $\overline{G}$ runs over all adelic quotient bundles of $\overline{E}$.
We refer the readers to [9, §4.3] for more details and properties of these invariants.
2.2 Harder-Narasimhan filtration from $\mathbb{R}$-filtration
In this part, we introduce a method to filtrate an adelic vector bundle by its minimal slopes, which is called the Harder-Narasimhan filtration of an adelic vector bundle. We will use the language of $\mathbb{R}$-filtration to formulate this construction. See [5] for its origin and [9] for the version over adelic curve.
2.2.1 $\mathbb{R}$-filtrations on vector spaces
Let $K$ be a field, and $W$ be a $K$-vector space of finite dimension. For an $\mathbb{R}$-filtration on $W$, we mean a sequence $\mathcal{F}=\left(\mathcal{F}^{t}W\right)_{t\in\mathbb{R}}$ of vector subspaces of $W$, which satisfies the following conditions:
  1. (1)
    if $t\leqslant s$, then $\mathcal{F}^{s}W\subseteq\mathcal{F}^{t}W$;
  2. (2)
    $\mathcal{F}^{t}W=0$ for sufficiently large $t\in\mathbb{R}$, and $\mathcal{F}^{t}W=W$ for sufficiently small $t\in\mathbb{R}$;
  3. (3)
    the function
    $\begin{array}[]{rcl}\mathbb{R}&\longrightarrow&\mathbb{R}\\ t&\mapsto&\dim_{K}(\mathcal{F}^{t}W)\end{array}$
    is left continuous with respect to Euclidean topology.
2.2.2 A measure induced by $\mathbb{R}$-filtration
Let $K$ be a field. For a $K$-vector space $W$ of finite dimension equipped with an $\mathbb{R}$-filtration $\mathcal{F}$, we denote by $\nu_{(W,\mathcal{F})}$ (or by $\nu_{W}$ if there is no ambiguity) the Borel probability measure on $\mathbb{R}$, which is induced from the derivative of the function
$\begin{array}[]{rcl}\mathbb{R}&\longrightarrow&\mathbb{R}\\ t&\mapsto&-\frac{\dim(\mathcal{F}^{t}W)}{\dim(W)}.\end{array}$
In fact, $\nu_{(W,\mathcal{F})}$ is a linear combination of some Dirac measures. More precisely, if the $\mathbb{R}$-filtration $\mathcal{F}$ consists of a flag
$W=W_{0}\supsetneq W_{1}\supsetneq\cdots\supsetneq W_{n}=0$
together with the jumps at the real numbers $\lambda_{1}<\cdots<\lambda_{n}$, then we have
$\nu_{(W,\mathcal{F})}=\sum_{i=1}^{n}\frac{\dim\left(W_{i-1}/W_{i}\right)}{\dim(W)}\delta_{\lambda_{i}},$
where $\delta_{\lambda_{i}}$ is the Dirac operator at the real number $\lambda_{i}$.
Next, we define the function
$\begin{array}[]{rrcl}\lambda:&W&\longrightarrow&\mathbb{R}\cup\{+\infty\}\\ &x&\mapsto&\sup\{t\in\mathbb{R}\mid x\in\mathcal{F}^{t}W\}.\end{array}$ (1)
Then we have $\lambda(W)\subset\operatorname{supp}(\nu_{W})\cup\{+\infty\}$, and $\lambda(x)$ is finite if $x\neq 0$.
In addition, we define
$\lambda_{+}(W)=\int_{0}^{+\infty}x\nu_{W}(\mathrm{d}x)\text{, and }\lambda_{\max}(W)=\max\limits_{x\in W\smallsetminus\{0\}}\lambda(x).$ (2)
2.2.3 Harder-Narasimhan filtration
By the notion of $\mathbb{R}$-filtration, we introduce a filtration by the minimal slopes of adelic bundles. This is original from an analogical study of vector bundles on a smooth projective curve [13].
[…]
Let $\overline{E}$ be a non-zero adelic vector bundle over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. For any $t\in\mathbb{R}$, let
$\mathcal{F}_{\operatorname{HN}}^{t}\overline{E}=\bigcap_{\epsilon>0}\sum_{\begin{subarray}{c}\{0\}\neq F\in\Theta(E)\\ \widehat{\mu}_{\min}(\overline{F})\geqslant t-\epsilon\end{subarray}}F=\sum_{\begin{subarray}{c}\{0\}\neq F\in\Theta(E)\\ \widehat{\mu}_{\min}(\overline{F})\geqslant t\end{subarray}}F,$
where $\Theta(E)$ denotes the set of all vector subspaces of $E$, and $\overline{F}$ is equipped with the subspace norms. This $\mathbb{R}$-filtration $\mathcal{F}_{\operatorname{HN}}$ is called the Harder-Narasimhan filtration of $\overline{E}$.
The last equality in Definition 2.2.3 if from [9, Proposition 4.3.36]. In fact, the Harder-Narasimhan filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
$\{0\}=E_{1}\subsetneq E_{2}\subsetneq\cdots\subsetneq E_{n}=E$ (3)
and a decreasing sequence of real numbers $\mu_{1}>\cdots>\mu_{n}$ represents the jumps of the $\mathbb{R}$-filtration $(\mathcal{F}_{\operatorname{HN}}^{t}\overline{E})_{t\in\mathbb{R}}$, where
$\mu_{i}=\sup\left\{t\in\mathbb{R}\mid\dim_{K}(\mathcal{F}^{t}_{\operatorname{HN}}\overline{E})\geqslant i\right\}.$
In particular, still by [9, Proposition 4.3.36], we have $\mu_{n}=\widehat{\mu}_{\min}(\overline{E})$. The uniqueness of the flag (3) is from [9, Theorem 4.3.58].
2.3 Filtrations on graded algebras
In this part, we study some asymptotic properties of graded linear algebras by $\mathbb{R}$-filtraion.
2.3.1 Approximable graded algebras
The notation of approximable graded algebra is from [4], whose idea is original from [18].
[…]
Let $K$ be a field, and $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be a graded $K$-algebra. We define that $E_{\bullet}$ is approximable if:
  1. (1)
    for each $D\in\mathbb{N}$, $E_{D}$ is of finite dimension, and $E_{D}\neq 0$ for sufficiently large $D\in\mathbb{N}$;
  2. (2)
    for every $\epsilon\in(0,1)$, there exists a $p_{0}\in\mathbb{N}^{+}$, such that for any $p\geqslant p_{0}$, we have
    $\liminf\limits_{n\to\infty}\frac{\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{p}\rightarrow E_{np}\right)\right)}{\dim_{K}(E_{np})}>1-\epsilon,$
    where $S^{n}B_{p}\rightarrow E_{np}$ is the canonical homomorphism defined by the algebra structure on $E_{\bullet}$.
2.3.2 Convergence of measures on an approximable algebra
For a graded algebra, we equip an $\mathbb{R}$-filtration with its each homogeneous part. If it is an approximable graded algebras, we will have some convergent properties.
Let $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be an approximable graded algebra. Suppose for every $D\geqslant 0$, the vector space $E_{D}$ is equipped with an $\mathbb{R}$-filtration $\mathcal{F}$. We ignore the index $D\in\mathbb{N}$ since there will be no ambiguity.
[…]
With all the above notations and conditions. If for every homogeneous elements $x_{1},x_{2}\in E_{\bullet}$, we have
$\lambda(x_{1}x_{2})\geqslant\lambda(x_{1})+\lambda(x_{2}),$
where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees of $x_{1}$ and $x_{2}$ are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
For every $\epsilon>0$, let $T_{\epsilon}$ be the operator on the space of Borel probability measures on $\mathbb{R}$ by mapping $x\mapsto\epsilon x$.
We have the following result of filtrated graded algebras.
[…]
With all the above notations. Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated approximable graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
$\sup_{n\geqslant 1}\frac{\lambda_{\max}(E_{n})}{n}<+\infty,$
where $\lambda_{\max}(\raisebox{1.72218pt}{.})$ is defined at (2). Then the sequence $\left(\frac{\lambda_{\max}(E_{n})}{n}\right)_{n\in\mathbb{N}^{+}}$ converges in $\mathbb{R}$, and the sequence of measures $\left(T_{\frac{1}{n}}(\nu_{E_{n}})\right)_{n\geqslant 1}$ converges weakly to a Borel probability measure on $\mathbb{R}$.
2.3.3 Limits of a measure from filtrated algebras
We consider the filtrated approximable graded $K$-algebra $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$. Let $\nu_{E_{\bullet}}$ be the limit of the sequence of measures $\left(T_{\frac{1}{n}}(\nu_{E_{n}})\right)_{n\geqslant 1}$, and $\lambda^{\operatorname{asy}}_{\max}(E_{\bullet})=\lim\limits_{n\to\infty}\frac{\lambda_{\max}(E_{n})}{n}$, whose existences are assured by Proposition 2.3.2. In addition, we have
$\lambda_{+}^{\operatorname{asy}}(E_{\bullet}):=\int_{0}^{+\infty}x\nu_{E_{\bullet}}(\mathrm{d}x)=\lim\limits_{n\to\infty}\frac{\lambda_{+}(E_{n})}{n},$ (4)
where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [4, §2.4] for more details.
We have the following result.
[…]
Let $E_{\bullet}$ be a filtrated approximable graded $K$-algebra. Then $\lambda_{+}^{\operatorname{asy}}(E_{\bullet})>0$ if and only if $\lambda_{\max}^{\operatorname{asy}}(E_{\bullet})>0$.
3 Graded linear series of global section spaces
Let $S$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, and $\overline{L}$ be an adelic line bundle on $X$ over $S$. For our purpose, we will construct some graded linear series consisting of some particular adelic vector bundles over $S$ from above data.
3.1 Adelic line bundles on an arithmetic variety
First, we recall some notions about arithmetic positivities of the adelic line bundles on arithmetic variety over an adelic curve, which are original from [8, 11]. We refer the readers to [11, §6 – §9] for a self-contained introduction to these notions.
3.1.1 Adelic line bundle
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $X$ be a projective scheme over $\operatorname{Spec}K$. We say that $\overline{L}=(L,\varphi)$ is an adelic line bundle on $X$, if $L$ is a line bundle on $X$, and $\varphi=(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is a family of metrics on $L_{w}$ for each $w\in\Omega$, which is dominated and measurable.
In addition, we suppose that either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Then we construct a family of norms on $H^{0}(X,L)$ in the way below: for every $w\in\Omega$ and $x\in X_{w}^{\operatorname{an}}$, define
|.$\begin{array}[]{rrcl}|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x):&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&|s|_{\varphi_{w}}(x)\end{array}$
as the value of the global section $s$ at the point $x$ with respect to the given metric, and
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}}:&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&\sup\limits_{x\in X^{\operatorname{an}}_{w}}\{|s|_{\varphi_{w}}(x)\}.\end{array}$ (5)
We denote $\varphi=(\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}})_{w\in\Omega}$ the norm family on $H^{0}(X,L)$ over $S$ if there is no ambiguity. By [9, Theorem 6.1.13, Theorem 6.1.32], $(H^{0}(X,L),\varphi)$ is an adelic vector bundle over $S$.
Let $\overline{L}=(L,\varphi)$ and $\overline{M}=(M,\psi)$ be two adelic line bundles on $X$. We denote by $\overline{L}\otimes\overline{M}=(L\otimes M,\varphi+\psi)$ the tensor adelic line bundle, where $\varphi+\psi$ is the tensor product of metrics. In particular, for $n\in\mathbb{N}$, we denote by $\overline{L}^{\otimes n}=(L^{\otimes n},n\varphi)$ the adelic line bundle of $n$-times tensor product of $\overline{L}$, and by $(H^{0}(X,L^{\otimes n}),n\varphi)$ the adelic vector bundle over $S$ defined above. By [9, Proposition 6.1.28, Proposition 6.1.31], all the above constructions are all well-defined.
We refer the readers to [9, §6.1] for a precise explanation of §3.1.1, see also [11, §2.6 – §2.8].
3.1.2 Semi-positive metric
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, $L$ be a line bundle on $X$, $\varphi=(\varphi_{w})_{w\in\Omega}$ and $\psi=(\psi_{w})_{w\in\Omega}$ be two families of metrics on $L$ such that $(L,\varphi)$ and $(L,\psi)$ are both adelic line bundles. Then we define the distance between $\varphi$ and $\psi$ as
$d(\varphi,\psi)=\int_{\Omega}\sup_{x\in X^{\operatorname{an}}_{w}}\left|\log\frac{|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x)}{|\raisebox{1.72218pt}{.}|_{\psi_{w}}(x)}\right|\nu(\mathrm{d}w).$
If $L$ is semi-ample, and if there exist a positive integer $m$ and a sequence $(\varphi_{n})_{n\in\mathbb{N}}$ of quotient metric families (where $\varphi_{n}$ a metric family of $L^{\otimes mn}$), such that
$\lim_{n\to\infty}\frac{1}{mn}d(nm\varphi,\varphi_{n})=0,$
we say that the metric family $\varphi$ is semi-positive.
3.1.3 Ampleness
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. We say $\overline{L}$ is relatively ample if $L$ is ample and $\varphi$ is semi-positive.
In addition, an adelic line bundle $\overline{L}$ on $X$ is ample, if it is relatively ample and if there exists an $\epsilon>0$ such that the inequality
$(\overline{L}|_{Y}^{\dim(Y)+1})\geqslant\epsilon\deg_{L}(Y)(\dim(Y)+1)$
holds for any integral closed subscheme $Y$ of $X$, where $(\overline{L}|_{Y}^{\dim(Y)+1})$ is the arithmetic intersection number studied in [10]. We also call it arithmetically ample to distinguish it from the geometric case.
[…]
We compare the above definition with the arithmetically ample definition of S. Zhang in [29], where an arithmetically ample Hermitian line bundle on an arithmetic variety requires the metrics on the line bundle to be smooth. The definition in [11, Definition 9.1.1] over an adelic curve only require continuous metrics. In this article, we follow the definition given in [11, Definition 9.1.1] as above.
3.1.4 Nefness
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. We say that $\overline{L}$ is relatively nef if there exists a relatively ample adelic line bundle $\overline{A}$ on $X$ and a positive integer $N$ such that, for any $n\geqslant N$, the tensor product of adelic line bundles $\overline{L}^{\otimes n}\otimes\overline{A}$ is relatively ample.
An adelic line bundle $\overline{L}$ is said to be nef or arithmetically nef, if there exist an arithmetically ample line bundle $\overline{A}$ and a positive integer $M$ such that $\overline{L}^{\otimes m}\otimes\overline{A}$ is ample for any $m\geqslant M$ in the arithmetic sense.
3.1.5 Effectivity
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. If $s\in H^{0}(X,L)$ is a non-zero global section which satisfies $\|s\|_{\varphi_{w}}\leqslant 1$ for all $w\in\Omega$, we say that the global section $s$ is effective.
We say that an adelic line bundle $\overline{L}$ is effective or arithmetically effective, if it admits at least one effective global section by the above definition.
3.2 Algebraic construction of graded linear series
In this part we will only consider the geometric setting, which means everything works over an arbitrary field with no metric. For every $D\in\mathbb{N}^{+}$, we consider the global section space $H^{0}(X,L^{\otimes D})$, and the graded $K$-algebra $\bigoplus\limits_{n\geqslant 0}H^{0}(X,L^{\otimes n})$.
[…]
Let $X$ be a projective variety over $K$, and $L$ be a line bundle on $X$. For a graded linear series of $L$, we mean a graded sub-$K$-algebra of $\bigoplus\limits_{n\geqslant 0}H^{0}(X,L^{\otimes n})$.
3.2.1 Graded linear series containing an ample divisor
The following definition is original from [18, Definition 2.9], where it is called "Condition (C)". This notion plays a key role in the arithmetic Fujita approximation of [4], and we follow the notation of [4, Definition 3.1].
[…]
Let $X$ be a projective variety, and $L$ be a line bundle on $X$. Suppose that there exists a $p\in\mathbb{N}^{+}$, an ample line bundle $A$, an effective line bundle $M$ on $X$, and a non-zero section $s\in H^{0}(X,M)$, such that $L^{\otimes p}=A\otimes M$. In this case, we say a graded linear series $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ of $L$ contains an ample divisor if the homomorphism of graded algebras
$\bigoplus_{n\geqslant 0}H^{0}(X,A^{\otimes n})\rightarrow\bigoplus_{n\geqslant 0}H^{0}(X,L^{\otimes np})$
induced by $s$ can factor through $\bigoplus\limits_{n\geqslant 0}W_{np}$.
In §2.3.2, we have seen that a filtrated approximable graded linear series has many convergent properties. The following property is original from [18], which is applied in [4] to judge whether a graded linear series is approximable.
[…]
If $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ contains an ample divisor and $W_{n}\neq 0$ for sufficiently large $n$, then $W_{\bullet}$ is approximable.
3.2.2 Big line bundles
Let $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ be a graded linear series of $L$. We denote
$\operatorname{vol}(W_{\bullet})=\limsup_{n\to\infty}\frac{\dim_{K}(W_{n})}{n^{\dim(X)}/\dim(X)!}.$
If $W_{n}=H^{0}(X,L^{\otimes n})$, we denote $\operatorname{vol}(L)=\operatorname{vol}(W_{\bullet})$.
If $\operatorname{vol}(L)>0$, we say that $L$ is a big line bundle on $X$.
3.3 Arithmetic invariants of graded linear series
In this part, we consider some arithmetic constructions.
3.3.1
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity. By [9, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$. Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
For every $D\in\mathbb{N}$, let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$. In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
3.3.2
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over $S$. We define the arithmetic volume of $\overline{E}_{\bullet}$ as
$\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\limsup_{n\to\infty}\frac{\widehat{\deg}_{+}\left(\overline{E}_{n}\right)}{n^{\dim(X)+1}/(\dim(X)+1)!}.$ (6)
In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicity.
If $\widehat{\operatorname{vol}}(\overline{L})>0$, we say that $\overline{L}$ is big on $X$, or arithmetically big to distinguish it from the geometric case.
3.3.3
For every $D\in\mathbb{N}$, let $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E}_{D})\right)_{t\in\mathbb{R}}$ be the $\mathbb{R}$-filtration induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ following Definition 2.2.3, where we ignore the homogeneous degree $D\in\mathbb{N}$ of $\overline{E}_{D}$.
We denote by $\nu_{\overline{E}_{D}}$ the Borel probability measure induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ constructed in the sense of §2.2.2. Then we have the result below from Proposition 4.3.50, Proposition 4.3.51 and Corollary 4.3.52 of [9], see also [11, §2.5.4].
[…]
With all the above notations and conditions. We have
$0\leqslant\widehat{\deg}_{+}(\overline{E}_{D})-\int_{0}^{+\infty}x\nu_{\overline{E}_{D}}(\mathrm{d}x)\leqslant\frac{1}{2}\nu(\Omega_{\infty})\dim_{K}(E_{D})\log(\dim_{K}(E_{D})).$
3.4 Graded linear series over trivially valued fields
Let $K$ be a field, $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a sub-field which is dense in $K_{w}$ for all $w\in\Omega$, and $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field introduced in Example 2.1.1.
In fact, for any normed vector space over $(K,|\raisebox{1.72218pt}{.}|_{0})$ of finite dimension, its norm takes only finitely many values (cf. [11, §3.1]).
In this part, we will study the adelic vector bundles and graded linear series over $S_{0}$ following that in [11, §3]. Some constructions are from those over $S$.
3.4.1
Let $\overline{E}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{E}$ defined in Definition 2.2.3. We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $E$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&E&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\}\right).\end{array}$ (7)
Then $\overline{E}_{\operatorname{HN}}=\left(E,\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}\right)$ is an adelic vector bundle over $S_{0}$.
3.4.2
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$ indexed by $D\in\mathbb{N}$. Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an adelic vector bundle over $S_{0}$.
By [11, Proposition 4.1.3], for every $s\in E_{m}$ and $t\in E_{n}$, we have
$\|s\cdot t\|^{\operatorname{HN}}_{m+n}\leqslant\exp(f(m)+f(n))\|s\|^{\operatorname{HN}}_{m}\|t\|^{\operatorname{HN}}_{n},$
where $f(n)=\frac{3}{2}\nu(\Omega_{\infty})\log n$, and $\nu(\Omega_{\infty})$ is the measure of archimedean absolute values of $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. The term $f(n)$ is original from the partial response of the semi-stable conjecture of Bost [1, Theorem A], see also [8, §5.6].
In order to obtain a new family of norms which are sub-multiplicative, we introduce the spectral norm .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ of .$\|\raisebox{1.72218pt}{.}\|^{\operatorname{HN}}_{D}$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{nD}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
In this case, $\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}\right)$ is an adelic vector bundle over $S_{0}$. By [11, Proposition 3.2.2 (3)], for every $s\in E_{m}$ and $t\in E_{n}$, we have
$\|s\cdot t\|_{\mathrm{sp},m+n}\leqslant\|s\|_{\mathrm{sp},m}\|t\|_{\mathrm{sp},n}.$
3.4.3
Let $\overline{E}_{\operatorname{HN},D}=\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}\right)$, then
$\overline{E}_{\operatorname{HN},\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{\operatorname{HN},D}$
is a filtrated graded linear series (see Definition 2.3.2 for the definition) over $S_{0}$, where we take $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ in Definition 2.3.2.
We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we omit the index $D\in\mathbb{N}$ in the above notation. The existence of such $\mathcal{F}_{\mathrm{sp}}$ is ensured by the finiteness of the values of .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ on $\overline{E}_{\operatorname{HN},D}$.
3.4.4
Some asymptotic behaviors of $\overline{E}_{\operatorname{HN},\bullet}$ are useful for the study of $\overline{E}_{\bullet}$. Since $\overline{E}_{\operatorname{HN},\bullet}$ is filtrated, Proposition 2.3.2 and Proposition 2.3.3 will be applied in this process.
[…]
With all the notations in §3.4.2 and §3.4.3, we have
$\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\widehat{\operatorname{vol}}(\overline{E}_{\operatorname{HN},\bullet}).$
Proof.
It is from [11, Corollary 3.4.4, (4.1.2), Corollary-Definition 4.1.4] directly. ∎
By Proposition 2.3.2, the sequence $\left(\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\right)_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$. If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the superior norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
The following result is an equivalent property of arithmetically bigness.
[…]
Let $\overline{L}$ be an adelic line bundle on $X$ over a proper adelic curve $S$. If $L$ is a big line bundle on $X$, then $\widehat{\operatorname{vol}}(\overline{L})>0$ if and only if $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$. In addition, we have
$\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})=\lambda_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}).$
Proof.
The equivalence of $\widehat{\operatorname{vol}}(\overline{L})>0$ and $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$ is from [9, Proposition 6.4.18] directly. The equality is from [11, Proposition 3.4.1]. ∎
From Proposition 3.3.3 and Proposition 3.4.4, we have the following result by an elementary calculation.
[…]
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$ be a normed graded linear series over $S$. If $\operatorname{vol}(E_{\bullet})>0$ following the definition in §3.2.2, then we have
$\lambda_{+}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})=\frac{\widehat{\operatorname{vol}}(\overline{E}_{\bullet})}{\left(\dim(X)+1\right)\operatorname{vol}(E_{\bullet})},$
where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure introduced in §3.4.3 over $S_{0}$.
3.4.5
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over an arbitrary proper curve $S$. We consider $\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)$, and we define
$\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})=\liminf_{n\to\infty}\frac{\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)}{n},$
where the adelic vector bundle $(H^{0}(X,L^{\otimes n}),n\varphi)$ over $S$ is defined in §3.1.1.
We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [11, §9] below, where we always suppose that $K$ is perfect.
[…]
Let $\overline{L}$ be a relatively ample adelic line bundle. Then the following properties are equivalent:
  1. (1)
    $\overline{L}$ is ample;
  2. (2)
    $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>0$;
  3. (3)
    there exists an $\epsilon>0$, such that for any closed sub-scheme $Y$ of $X$, we have $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}|_{Y})>\epsilon$.
See §3.1.3 for the definitions of relatively ample and ample adelic line bundles.
[…]
Let $\overline{L}$ be an adelic line bundle. Then the following properties are equivalent:
  1. (1)
    $\overline{L}$ is nef;
  2. (2)
    $\overline{L}$ is relatively nef and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})\geqslant 0$.
See §3.1.4 for the definitions of relatively nef and nef adelic line bundles.
4 Fujita approximation of graded linear series
In this section, we will always denote $E_{D}=H^{0}(X,L^{\otimes D})$ for simplicity, and all the constructions will be based on this notation.
4.1 Approximatible graded linear series
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable which is dense in $K_{w}$ for all $w\in\Omega$. Let $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field defined in Example 2.1.1. In this part, we will consider some approximated properties of normed graded linear series over $S_{0}$ constructed from that over $S$.
4.1.1
Let $\lambda\in\mathbb{R}_{\geqslant 0}$, $\overline{E}_{\operatorname{HN},0}^{\lambda}=K$, and
$\overline{E}_{\operatorname{HN},D}^{\lambda}=\operatorname{Vect}_{K}\left\{s\in H^{0}(X,L^{\otimes D})\mid\|s\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)\right\}$ (9)
for $D\geqslant 1$. So we have
$\overline{E}_{\operatorname{HN},D}^{\lambda}=\mathcal{F}_{\mathrm{sp}}^{D\lambda}\overline{E}_{\operatorname{HN},D},$
where the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §2.3.2, and $\overline{E}_{\operatorname{HN},D}$ follows the definition in §3.4.3 by taking $E_{D}=H^{0}(X,L^{\otimes D})$. In addition, we define
$\overline{E}^{\lambda}_{\operatorname{HN},\bullet}=\bigoplus_{D\geqslant 0}\overline{E}_{\operatorname{HN},D}^{\lambda},$ (10)
which is a graded linear series over $S_{0}$.
4.1.2
By the properties of $\overline{E}_{\operatorname{HN},\bullet}$ in §3.4.3, we have the following proposition.
[…]
The graded linear series (10) is a filtrated graded linear series over $S_{0}$.
4.1.3
Next, we prove that $\overline{E}_{\operatorname{HN},D}^{\lambda}$ is approximable when $\lambda$ is of some particular values. For this target, we will prove it contains an ample divisor, and see Definition 4.1.3.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S$. With all the notations and assumption same as in (10). If $\overline{L}$ is big, then for all $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda}_{\operatorname{HN},\bullet}$ contains an ample divisor.
Proof.
By Proposition 3.4.4 and Proposition 3.4.4, from the fact that $\overline{L}$ is big, we have $\overline{E}_{\operatorname{HN},\bullet}$ is big.
In addition, by the choice of $\lambda$, we have $\lambda^{\operatorname{asy}}_{\max}(\overline{E}_{\operatorname{HN},\bullet}^{\lambda})>0$ by definition directly. Then for a sufficiently large $n\in\mathbb{N}$, $L^{\otimes n}$ has a global section $s\in H^{0}(X,L^{\otimes n})$ satisfying $\|s\|_{\mathrm{sp},n}\leqslant\exp(-n\lambda)$, which proves $\overline{E}_{\operatorname{HN},n}^{\lambda}\neq 0$ for $n\in\mathbb{N}$ sufficiently large.
By Kodaira decomposition [16, Corollary 2.2.7], there exists a $p_{0}\in\mathbb{N}$, such that for every $p\in\mathbb{N}\cap[p_{0},+\infty)$, there always exist an ample line bundle $M$ and an effective line bundle $N$, which satisfies $L^{\otimes p}=M\otimes N$.
We may choose the above $p$ large enough, such that $\bigoplus\limits_{D\geqslant 0}H^{0}(X,M^{\otimes D})$ is generated by global sections $t_{1},\ldots,t_{l}\in H^{0}(X,M)$, which is also considered as sections in $H^{0}(X,L^{\otimes d})$.
We denote by $\lambda_{0}=\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity. Then for all $\epsilon>0$, there exist a $k\in\mathbb{N}$ and an $s\in H^{0}(X,L^{\otimes k})$, such that
$\|s\|_{\mathrm{sp},pk}<\exp\left(-(\lambda_{0}-\epsilon/2)k\right).$
Then there exists a $d\in\mathbb{N}$, such that
$\|t_{i}s^{d}\|_{\mathrm{sp},kd+p}<\exp\left(-(\lambda_{0}-\epsilon)(kd+p)\right),$
where $i=1,\ldots,l$.
Let $G_{\bullet}^{\epsilon}$ be the graded linear series generated by $s^{D}t_{1},\ldots,s^{D}t_{l}$, which is a sub-graded algebra of $E_{\operatorname{HN},\bullet}$. Then for $D\in\mathbb{N}$, we have
$G_{D}^{\epsilon}=\operatorname{Vect}_{K}\left\{(s^{d}t_{1})^{a_{1}}\cdots(s^{d}t_{l})^{a_{l}}\mid a_{1}+\cdots+a_{l}=D\right\}=s^{dD}H^{0}(X,M^{\otimes D}),$
which is a sub-$K$-algebra of $E_{\operatorname{HN},(kd+p)D}$. Then for all $\lambda\in(-\infty,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}))$, we embed $H^{0}(X,M^{\otimes D})$ into $E_{\operatorname{HN},pD}^{\lambda}$ via $s^{dD}$. Then the homomorphism
$\bigoplus_{D\geqslant 0}H^{0}(X,M^{\otimes D})\longrightarrow\bigoplus_{D\geqslant 0}H^{0}(X,L^{\otimes pD})$
factors through $\bigoplus\limits_{n\geqslant 0}E_{\operatorname{HN},np}^{\lambda}$. ∎
[…]
In [4, Proposition 3.11], a similar result of Proposition 4.1.3 is proved by an arithmetic analogue of Kodaira decomposition [25, Corollary 2.4]. Until the author’s knowledge, we do not know whether it is still valid over a general adelic curve. To avoid this obstruction, we use a method quite similar to that of [2], and see also [9, §6.3].
Same as the argument in [4, Corollary 3.13], we have the following result.
[…]
We keep all the notations and conditions in Proposition 4.1.3. The graded linear series (10) over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ is approximable for any $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$.
Proof.
By Proposition 3.2.1, we have the assertion from Proposition 4.1.3 directly. ∎
4.2 Fujita approximation of graded linear series
Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, where we follow the notation in (9). In this case, we define $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ to be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ defined in Example 2.1.1.
The following result is an analogue of [4, Theorem 4.1] over $S_{0}$, whose proof is quite similar to the ancient one.
[…]
Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. With all the above notations and constructions, we have
$\widehat{\operatorname{vol}}(\overline{L})=\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
Proof.
For any $n\in\mathbb{N}^{+}$, let $\nu_{\mathrm{sp},n}=T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$, where the notation $T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$ is introduced in §2.3.2, and the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §3.4.2 with respect to the spectral norms. Then by Proposition 2.3.2, the sequence $(\nu_{\mathrm{sp},n})_{n\geqslant 1}$ converges weakly to a Borel probability measure, which is denoted by $\nu_{\mathrm{sp}}$.
Similarly, let $\nu_{\mathrm{sp},n}^{(p)}=T_{\frac{1}{np}}\nu_{\left(E^{(p)}_{\operatorname{HN},np},\mathcal{F}_{\mathrm{sp}}\right)}$, which also converges weakly to a Borel probability measure noted by $\nu_{\mathrm{sp}}^{(p)}$ by Proposition 4.1.2.
In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ can be approximated well by $\nu_{\mathrm{sp}}^{(p)}$ for $p\in\mathbb{N}$.
Step I. – Simultaneous approximations.
Let
[…]
be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}))$ such that
$\nu_{\mathrm{sp}}(\{t_{1},\ldots,t_{m}\})=0.$ (11)
We define the function
$\begin{array}[]{rrcl}h_{D}:&\mathbb{R}&\longrightarrow&\mathbb{R}\\ &x&\mapsto&\sum\limits_{i=0}^{m-1}t_{i}\mathbbm{1}_{[t_{i},t_{i+1})}(x)+t_{m}\mathbbm{1}_{[t_{m},+\infty)}(x).\end{array}$
By Corollary 4.1.3 and Definition 2.3.1, for every $\epsilon>0$, there exists a $p\in\mathbb{N}$ sufficiently large, such that $E_{\bullet}^{(p)}$ approximates all $E^{t_{i}}_{\operatorname{HN},\bullet}$ defined at (10) simultaneously, where $i=0,\ldots,m$. In other words, there exists an $N_{0}\in\mathbb{N}$, such that for every $n\geqslant N_{0}$, we have
$\inf_{0\leqslant i\leqslant m}\frac{\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{\operatorname{HN},p}^{t_{i}}\rightarrow E_{\operatorname{HN},np}^{t_{i}}\right)\right)}{\dim_{K}(E_{\operatorname{HN},np}^{t_{i}})}\geqslant 1-\epsilon$
for all $i=0,\ldots,m$. Then by the above inequality, we obtain
$\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{npt_{i}}(E_{\operatorname{HN},np}^{(p)}))$$\geqslant$ $\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{\operatorname{HN},p}^{t_{i}}\rightarrow E_{\operatorname{HN},np}^{t_{i}}\right)\right)$
$\geqslant$ $(1-\epsilon)\dim_{K}(E_{\operatorname{HN},np}^{t_{i}})$
for all $i=0,\ldots,m$.
Step II. – An upper bound of integrations.
We note that
$np\dim_{K}(E_{\operatorname{HN},np}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp},np}^{(p)}(\mathrm{d}t)=-\int_{0}^{+\infty}t\mathrm{d}\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{t}E_{\operatorname{HN},np}^{(p)})$
$=$ $\int_{0}^{+\infty}\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{t}E_{\operatorname{HN},np}^{(p)})\mathrm{d}t\geqslant\sum_{i=1}^{m}np(t_{i}-t_{i-1})\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{npt_{i}}E_{\operatorname{HN},np}^{(p)})$
$\geqslant$ $np(1-\epsilon)\sum_{i=1}^{m}(t_{i}-t_{i-1})\dim_{K}(E_{\operatorname{HN},np}^{t_{i}}),$
where the first inequality is from the decreasing of $\mathcal{F}_{\mathrm{sp}}$, and the last inequality is from (4.2) combined with the fact $E_{\operatorname{HN},np}^{t_{i}}=\mathcal{F}_{\mathrm{sp}}^{npt_{i}}E_{\operatorname{HN},np}$.
In addition, by applying Abel’s summation formula to the last term of the above inequality, we have
$\dim_{K}(E_{\operatorname{HN},np}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp},np}^{(p)}(\mathrm{d}t)\geqslant(1-\epsilon)\dim_{K}(E_{\operatorname{HN},np})\int_{0}^{+\infty}h_{D}(t)\nu_{\mathrm{sp},np}(\mathrm{d}t).$
Step III. – Estimates of arithmetic volumes.
Let $d=\dim(X)$. By Proposition 3.4.4, we have
$\lim_{n\to\infty}\frac{d+1}{(np)^{d+1}/(d+1)!}\dim_{K}(E_{\operatorname{HN},np}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp},np}^{(p)}(\mathrm{d}t)$
$=$ $(d+1)\operatorname{vol}(E_{\bullet}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp}}^{(p)}(\mathrm{d}t)=\widehat{\operatorname{vol}}(E_{\bullet}^{(p)}).$
Then we have
$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)})$$\geqslant$ $\lim_{n\to\infty}\frac{d+1}{(np)^{d+1}/(d+1)!}(1-\epsilon)\dim_{K}(E_{\operatorname{HN},np})\int_{0}^{+\infty}h_{D}\mathrm{d}\nu_{\mathrm{sp},np}$
$=$ $(d+1)(1-\epsilon)\operatorname{vol}(L)\int_{\mathbb{R}}h_{D}\mathrm{d}\nu_{\mathrm{sp}},$
where the last equality is from [3, Chap. IV §5 $n^{\circ}$ 12 Proposition 22].
We choose a sequence of subdivisions $(D_{j})_{j\in\mathbb{N}}$ satisfying (11), such that the sequence of functions $(h_{D_{j}}(t))_{j\in\mathbb{N}}$ converges uniformly to
$\max\{t,0\}-\max\{t-\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}),0\}$
when $j\to\infty$. By definition, the support of $\nu_{\mathrm{sp}}$ is bounded from above by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})$, we obtain
$\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(E_{\bullet}^{(p)})\geqslant(d+1)(1-\epsilon)\operatorname{vol}(L)\int_{0}^{+\infty}t\nu_{\mathrm{sp}}(\mathrm{d}t)=(1-\epsilon)\widehat{\operatorname{vol}}(\overline{L}),$
where the last equality is from Proposition 3.4.4 on the equality of arithmetic volumes over $S$ and $S_{0}$. ∎
[…]
For a general graded linear series, if it satisfies the property of Proposition 4.2, we say that this graded linear series satisfies Fujita approximation. We would like to emphasize that this notion is different from those in [7, Definition 3.7] or [8, §6], though they are somewhat similar.
5 Fujita approximation of adelic line bundles
From the Fujita approximation of graded linear series proved in Proposition 4.2, we will prove the arithmetic Fujita approximation of an arithmetically big line bundle with coefficients in $\mathbb{Z}$, $\mathbb{Q}$ and $\mathbb{R}$.
5.1 Fujita approximation of big adelic line bundles
First, we will prove the arithmetic Fujita approximation of usual adelic line bundles. We will prove the nef version first, and apply it to prove the ample version. In order to apply Proposition 3.4.5, we assume the base field is perfect.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big line bundle on $X$ over $S$. Then for every $\epsilon>0$, there exists a birational morphism $v:X^{\prime}\rightarrow X$ together with a positive integer $p$, such that the decomposition $v^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
  1. (1)
    $\overline{M}$ is effective and $\overline{G}$ is arithmetically nef;
  2. (2)
    $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
Proof.
We devide the proof into three steps.
Step I. – Determining the geometric decomposition.
Let $\pi:X\rightarrow\operatorname{Spec}K$ be the structural morphism. For any $p\geqslant 1$ such that $E_{\operatorname{HN},p}^{0}\neq 0$ (see Proposition 4.1.3 for the existence of such $p\in\mathbb{N}$), let
$\phi_{p}:X_{p}\rightarrow X$
be the blow-up along the base locus of $E_{\operatorname{HN},p}^{0}$, which means
$X_{p}=\operatorname{Proj}\left(\operatorname{Im}\left(\bigoplus_{n\geqslant 0}\pi^{*}(E_{\operatorname{HN},np}^{(p)})\longrightarrow\bigoplus_{n\geqslant 0}L^{\otimes np}\right)\right).$
Let $G_{p}=\mathcal{O}_{X_{p}}(1)$, and $M_{p}$ be the line bundle on $X_{p}$ with respect to the exceptional divisor. Let $s$ be a global section of $M_{p}$ which trivializes $M_{p}$ outside the exceptional divisor. Then by definition, we have
$\phi_{p}^{*}(L^{\otimes p})=G_{p}\otimes M_{p}.$
The canonical morphism
$\phi_{p}^{*}\pi^{*}(E_{\operatorname{HN},p}^{0})\rightarrow G_{p}$
induces a projective embedding $i_{p}:X_{p}\rightarrow\mathbb{P}\left(E_{\operatorname{HN},p}^{0}\right)$ with $i^{*}\left(\mathcal{O}_{E_{\operatorname{HN},p}^{0}}(1)\right)=G_{p}$. Then the restriction of global sections of $\mathcal{O}_{E_{\operatorname{HN},p}^{0}}(n)$ on $X_{p}$ gives an injective homomorphism
$\operatorname{Im}\left(S^{n}E_{\operatorname{HN},p}^{0}\rightarrow E_{\operatorname{HN},np}\right)=E_{\operatorname{HN},np}^{(p)}\rightarrow H^{0}(X_{p},G_{p}^{\otimes n}),$ (13)
where $H^{0}(X_{p},G_{p}^{\otimes n})$ is considered as a sub-$K$-vector space of $H^{0}(X_{p},\phi_{p}^{*}(L^{\otimes p}))$ via the section $s$.
Step II. – Constructing the metrics of line bundles.
We consider $\overline{E}^{(p)}_{\bullet}$ as a normed graded linear series over $S$, which is equipped with the subspace norms of $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$. For any $n\geqslant 1$ and $w\in\Omega$, let .$\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
$\phi_{p}^{*}\pi^{*}(E_{\operatorname{HN},np}^{(p)})\rightarrow G_{p}^{\otimes n},$
where $E_{\operatorname{HN},np}^{(p)}$ is equipped with the superior norm .$\|\raisebox{1.72218pt}{.}\|_{\sup,w}$ with respect to $w\in\Omega$.
In fact, for any $x\in X_{p,w}(\overline{K}_{w})$ outside the exceptional divisor, it corresponds to a $1$-dimensional quotient of $E_{\operatorname{HN},p}^{0}$. For any $n\geqslant 1$, this quotient induces a $1$-dimensional quotient $\ell_{n,x}$ of $E_{\operatorname{HN},np}^{(p)}$.
By Hahn-Banach theorem, every linear form on $\ker(E_{\operatorname{HN},np}^{(p)}\rightarrow\ell_{n,x})$ can be extended to a linear form on $E_{\operatorname{HN},np}^{(p)}$. In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$, whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$. As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where .$\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$. Then for any section $u\in G_{p,w}$ over a neighborhood of $x$, we may assume that $u(x)^{\otimes n}$ equals to the image of $w(x)$ in $G_{p,w}^{\otimes n}(x)$ by dilation, then we have
$\|u(x)\|_{n,w}=\|w\|_{\sup,w}^{1/n}\geqslant\|w(x)\|_{w}^{1/n}=\|u(x)\otimes s(x)\|_{w}.$
Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ be a family of adelic metrics on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$. In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective adelic line bundle.
Step III. – The relation among arithmetic volumes.
By the construction in Step II, for any $n\geqslant 1$, we have
$p^{-(d+1)}\widehat{\operatorname{vol}}(G_{p},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n}),$
where
$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})=\lim_{m\rightarrow\infty}\frac{\widehat{\deg}_{+}(E_{\operatorname{HN},mp}^{(p)},(\|\raisebox{1.72218pt}{.}\|_{n,\sup,w})_{w\in\Omega})}{m^{d+1}/(d+1)!},$
and .$\|\raisebox{1.72218pt}{.}\|_{n,\sup,w}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [4, Corollary A.2], we have
$\limsup_{n\to\infty}\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
By Proposition 4.2, for any $\epsilon>0$, there exists a $p\in\mathbb{N}$, such that
$\limsup_{n\to\infty}p^{-(d+1)}\widehat{\operatorname{vol}}(G_{p},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\frac{\epsilon}{2}.$
Then there exists an $N_{0}\in\mathbb{N}$, such that
$p^{-(d+1)}\widehat{\operatorname{vol}}(G_{p},\alpha_{n})-\frac{\epsilon}{2}\geqslant\widehat{\operatorname{vol}}(\overline{L})-\frac{\epsilon}{2}$
for infinitely many $n\geqslant N_{0}$.
In addition, since $\widehat{\mu}_{\min}(\overline{E}_{\operatorname{HN},np}^{(p)})\geqslant 0$ for all $n\geqslant 1$, then it has a positive asymptotic minimal slope. Then by Proposition 3.4.5, we obtain that $(G_{p},\alpha_{n})$ is arithmetically nef with the fact that $K$ is perfect.
To sum up, we obtain the assertion. ∎
Next, we obtain the following ample version from Theorem 5.1 directly. The idea is absorbed from [4, Remark 4.4] and [11, Corollary 8.5.7].
[…]
We keep all the notations and conditions in Theorem 5.1. Then we are able to require $\overline{G}$ to be arithmetically ample in the term (1) of Theorem 5.1.
Proof.
Let $\overline{B}$ be an arithmetically big line bundle on $X$. Then by [9, Corollary 6.4.10], for $m\in\mathbb{N}$ sufficiently large, we have $\overline{L}^{\otimes m}\otimes\overline{B}^{\vee}$ is arithmetically big. By Theorem 5.1, for any $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$, a $p\in\mathbb{N}$, and a decomposition
$\phi^{*}(\overline{L}^{\otimes mp}\otimes\overline{B}^{\otimes(-p)})=\overline{G}\otimes\overline{M}$
with $\overline{G}$ arithmetically nef and $\overline{M}$ effective, such that
$p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L}^{\otimes m}\otimes\overline{B}^{\vee})-\epsilon.$
Let $\overline{B}^{\prime}$ be an arithmetically ample line bundle on $X^{\prime}$. Since $\phi^{*}(\overline{B})$ is arithmetically big, there exists a $q\in\mathbb{N}$, such that $\phi^{*}(\overline{B}^{\otimes q})\otimes\overline{B}^{\prime\vee}$ is effective. We consider
$\phi^{*}(\overline{L}^{\otimes mpq})=(\overline{G}^{\otimes q}\otimes\overline{B}^{\prime})\otimes(\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}).$
In the above decomposition, $\overline{G}^{\prime}:=\overline{G}^{\otimes q}\otimes\overline{B}^{\prime}$ is arithmetically ample, $\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}$ is effective, and
$(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\prime})\geqslant(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\otimes q})\geqslant m^{-(d+1)}(\widehat{\operatorname{vol}}(\overline{L}^{\otimes m}\otimes\overline{B}^{\vee})-\epsilon).$
By the continuity of the arithmetic volume function (see [9, Corollary 6.4.10]), we have the result. ∎
5.2 Case of $\mathbb{K}$-line bundles
In this part we will prove the arithmetic Fujita approximation is verified for adelic $\mathbb{Q}$-line bundles and $\mathbb{R}$-line bundles over adelic curves.
5.2.1
First, we prove the case of adelic $\mathbb{Q}$-line bundle, which is a quite straightforward corollary from the usual case.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big $\mathbb{Q}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
  1. (1)
    $\overline{M}$ and $\overline{G}$ are adelic line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
  2. (2)
    $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
Proof.
By definition, there exists an $m\in\mathbb{N}$, such that $\overline{L}^{\otimes m}$ is a usual arithmetically big line bundle. Then we obtain the assertion by applying Corollary 5.1 to $\overline{L}^{\otimes m}$. ∎
5.2.2
Next, we prove the case of adelic $\mathbb{R}$-line bundle, and see [9, §6.4.3] for an introduction to this notion. For this target, we prove an auxiliary lemma first, which allows us to approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle along an effective direction.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, and $X$ be a projective variety over $\operatorname{Spec}K$. Let $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$. Then for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and
$0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}(\overline{L}^{\prime})<\epsilon.$
Proof.
The adelic $\mathbb{R}$-line bundle $\overline{L}$ is arithmetically big, so $L$ is a big $\mathbb{R}$-line bundle. By an equivalent definition [16, Definition 2.2.21], we have $L=\sum\limits_{i=1}^{r}a_{i}L_{i}$, where $a_{1},\ldots,a_{r}>0$ and $L_{1},\ldots,L_{r}$ are usual big line bundles. By definition, $L_{1},\ldots,L_{r}$ are all effective in the geometric sense.
We equip metrics on $L_{1},\ldots,L_{r}$ over $S$, such that $\overline{L}_{1},\ldots,\overline{L}_{r}$ are arithmetically effective. Then for any $b_{1},\ldots,b_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{L}_{1}+\cdots+b_{r}\overline{L}_{r}$ is effective.
By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [9, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$, we always have
$0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}\left(\overline{L}-\sum_{i=1}^{r}t_{i}\overline{L}_{i}\right)<\epsilon.$
If there does not exist any arithmetically big $\mathbb{Q}$-line bundles satisfying the requirements of the lemma, then $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ can never be a $\mathbb{Q}$-line bundle for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$. It means there is no $\mathbb{Q}$-point in the region
$(a_{1}-\delta_{1},a_{1})\times\cdots\times(a_{r}-\delta_{r},a_{r})\subset\mathbb{R}^{r},$
which has a strict positive Lebesgue measure. Then it contradicts to the fact that $\mathbb{Q}^{r}$ is dense in $\mathbb{R}^{r}$ with respect to the Euclidean topology.
By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r}\in\mathbb{R}^{+}$ such that $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle. By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is effective. Let
$\overline{L}^{\prime}=\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i},$
which satisfies the requirement. ∎
We prove the Fujita approximation of $\mathbb{R}$-line bundles via Lemma 5.2.2.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
  1. (1)
    $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle, and $\overline{G}$ is an arithmetically ample line bundle;
  2. (2)
    $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
Proof.
By Lemma 5.2.2, for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and
$0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}(\overline{L}^{\prime})<\epsilon.$
Then we apply Corollary 5.2.1 to $\overline{L}^{\prime}$, and we obtain the assertion. ∎